# MD 20742, USA

### v2

@@ Appendix B Good targets for expanding maps
$\tau_{*}^{N}=\sum_{\overset{0<k_{1}<k_{2}<\dots<k_{r}\leq N}{{\operatorname{Sep}_{n}}(k_{1},\dots k_{r})=r}}\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N}}\right).$
Using $(M1)_{r},$ which has been proved, we have $\lim_{N\rightarrow\infty}\tau_{*}^{N}/\left(N\sigma_{\rho_{N}}\right)^{r}=1.$ Therefore, it suffices to show that
$\lim_{N\rightarrow\infty}\frac{\mu(A_{*}^{N})}{\tau_{*}^{N}}=1.$
By Boole’s inequality $\mu(A_{*}^{N})\leq\tau_{*}^{N}.$ For the lower bound we note that
@@
(B.1) $\mu\left(A_{*}^{N}\right)\geq\tau_{*,b}^{N}-\sum^{*}\mu(\Omega^{k_{1}}_{{\rho}_{N}}\cap\dots\cap\Omega^{k_{r}}_{{\rho}_{N}}\cap\Omega^{k_{r+1}}_{{\rho}_{N}}),$
where the sum in $\sum^{*}$ is taken over all tuples satisfying $\operatorname{Sep}(k_{1},\dots k_{r})=r.$ We now consider several cases depending on the size of $p=\min_{j=1,\dots,r}|k_{r+1}-k_{j}|$.
@@
(i) Since $x$ is non-periodic for each fixed $p_{0},$ we have that $\Omega_{{\rho}_{N}}\cap f^{-p}\Omega_{{\rho}_{N}}=\emptyset$ for all $p\leq p_{0}$ provided that $n$ is large enough. Hence the contribution of terms $p\leq p_{0}$ is zero.
@@
(ii) Next we suppose that $p_{0}<p\leq\varepsilon\operatorname{ln}N.$ Precisely we may assume that $k_{r+1}=k_{\bar{j}}+p$ for some ${\bar{j}}\leq r$ (the case where $k_{r+1}=k_{\bar{j}}-p$ is similar). Define $\bar{A}_{{\rho}_{N}}^{+}$ to be the characteristic function of $B\left(x,\left(\frac{a}{\rho_{N}-L}\right)^{s}\right).$ Then
$\bar{A}^{+}_{\rho_{N}}\geq{\mathbf{1}}_{\Omega_{\rho_{N}}},\,{\mathbb{E}}\bar{A}^{+}_{\rho_{N}}-\mu(\Omega_{\rho_{N}})\leq C\rho_{N}^{-(s+1)}\text{ and }\|\bar{A}^{+}_{\rho_{N}}\|_{BV}\leq 2\mu(\Omega_{\rho_{N}})+4.$
@@
${\hat{A}}_{j}=\begin{cases}\bar{A}_{{\rho}_{N}}^{+}&\text{if }j\neq{\bar{j}},\\ \bar{A}_{{\rho}_{N}}^{+}(\bar{A}_{{\rho}_{N}}^{+}\circ f^{p})&\text{if }j={\bar{j}}.\end{cases}$
If we denote by $M=\max_{x\in\mathbb{S}}|f^{\prime}(x)|,$ then
@@
$\leq$ $\left(\prod_{j\neq{\bar{j}}}\left({\mathbb{E}}{\hat{A}}_{j}+C\|\bar{A}_{{\rho}_{N}}^{+}\|_{BV}\theta^{K\operatorname{ln}N}\right)\right)\left({\mathbb{E}}{\hat{A}}_{\bar{j}}+C\|\bar{A}_{{\rho}_{N}}^{+}\|_{BV}^{2}M^{p}\theta^{K\operatorname{ln}N}\right)\leq\prod_{j=1}^{r}\left({\mathbb{E}}{\hat{A}}_{j}+n^{-10}\right)$
if $s(n)=K\operatorname{ln}N$ and $K$ is large enough. Noting that
${\mathbb{E}}({\hat{A}}_{\bar{j}})={\mathbb{E}}\left(\bar{A}_{{\rho}_{N}}^{+}(\bar{A}_{{\rho}_{N}}^{+}\circ f^{p})\right)\leq C\sigma_{{\rho}_{N}}\theta^{p}$
@@
by (2.4), we have that each term with $p$ as in the case (ii) is at most $C\sigma_{{\rho}_{N}}^{r}\theta^{p}.$ Summing over all $r$ tuples and over all $p\in(p_{0},\varepsilon\operatorname{ln}N]$ yields a total contribution of $CN^{r}\sigma_{\rho_{N}}^{r}\theta^{p_{0}}$ for some constant $C$ independent of $p_{0}.$
@@
(iii) The terms where ${\varepsilon}\operatorname{ln}N<p\leq K\operatorname{ln}N$ are $O(\sigma_{\rho_{N}}^{r+\eta})$ for some $\eta=\eta({\varepsilon})>0.$ Thus the total contribution of those terms are $O(N^{r}(\operatorname{ln}N)\;\sigma_{\rho_{N}}^{r+\eta}),$ which is also negligible.
@@
(iv) The terms where $p>K\operatorname{ln}N$ for sufficiently large $K$ is $O(\sigma_{\rho_{N}}^{r+1})$ similarly to the proof of $(M1)_{r}.$ Thus the total contribution of those terms are $O((N\sigma_{\rho_{N}})^{r+1})$ which is negligible compared to $\tau_{*}^{N}.$
Summing the estimates of cases (i)–(iii) above we see that the second term in the RHS of (B.1) is negligible comparing to the first, which proves $(M4)_{*}.$square
@@
Since $H_{N,{\bar{\varepsilon}}}\subset\tilde{H}_{k,{\bar{\varepsilon}}}$ for $2^{k}\leq N\leq 2^{k+1},$ it suffices to show that $\tilde{H}_{k,{\bar{\varepsilon}}}$ happens only for finitely many $k.$ Using (2.4) and similar method in the proof of Proposition 2.3 we conclude that
@@
Hence $\sum_{k}\mu(\tilde{H}_{k,{\bar{\varepsilon}}})<\infty$ and the result follows from the classical Borel-Cantelli Lemma. $\square$
@@
Let $p=\min_{1\leq j\leq m-1}\{k_{j+1}-k_{j}\}.$ Since $x$ is non-periodic, for each fixed $p_{0},$ we have that ${\hat{B}}_{N,j}\cap f^{-p}{\hat{B}}_{N,j}=\emptyset$ for all $p\leq p_{0}$ provided that $N$ is large enough because $x$ is non-periodic. Thus the contribution of the terms with tuples satisfying $p\leq p_{0}$ is zero for large $N.$
@@
We suppose that $\operatorname{Sep}(k_{1},\dots k_{m})=l<m.$ If $p_{0}<p\leq\varepsilon\operatorname{ln}N,$ Arguing as the case $(M4)_{*}$(ii) in the proof of Proposition 2.3, we conclude that each term is smaller than $C\mu({\hat{B}}_{N,j})^{l}\theta^{p}$ for fixed $p.$ Since the total number of terms is $O\left(N^{l}p^{m-l}\right),$ we have that total contribution of the terms with $p={\hat{p}}$ is $O\left(N^{l}{\hat{p}}^{m-l}\mu({\hat{B}}_{N,j})^{l}\theta^{\hat{p}}\right).$ The sum over tuples with ${\hat{p}}\in[p_{0},{\varepsilon}\operatorname{ln}N]$ gives a total contribution of $CN^{l}{p_{0}}^{m-l}\mu({\hat{B}}_{N,j})^{l}\theta^{p_{0}}$ for fixed $l$ and $p_{0}$ large enough (independent of $N$).
@@
On the other hand if $\varepsilon\operatorname{ln}N<p\leq K\operatorname{ln}N,$ then similarly to the proof of the case $(M4)_{*}$(iii) we obtain that
$\mu\left(\bigcap_{j=1}^{m}\left\{f^{n_{j}}y\in{\hat{B}}_{N,j}\right\}\right)\leq C\mu({\hat{B}}_{N,j})^{l+\eta}.$
@@
Summing over all $r$ tuples and over all $p\in(\varepsilon\operatorname{ln}N,K\operatorname{ln}N]$ yields a total contribution that is not greater than $C\left(N\mu({\hat{B}}_{N,j})\right)^{l}\mu({\hat{B}}_{N,j})^{\eta}\left(\operatorname{ln}N\right)^{m-l}.$

### v3

@@ Appendix B Good targets for expanding maps
$\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N}}\right)$ $\leq$ $C\left(\prod_{k_{j+1}-k_{j}>K\operatorname{ln}N}\sigma_{\rho_{N}}\right)(\theta^{\epsilon\operatorname{ln}N})^{r-m}\leq{\frac{C\sigma_{{\rho}_{N}}^{m}}{(\operatorname{ln}N)^{100r}}}.$
(c) Recall that
@@
$A_{*}^{N}$ $=$ $\left\{\omega\in\Omega:\,\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq N\,\,\text{such that}\,\,\operatorname{Sep}_{N}(k_{1},\dots,k_{r})=r\,\,\text{and}\,\,\omega\in\Omega^{k_{j}}_{{\rho}_{N}}\right\}$
$=$ $\bigcup_{\overset{0<k_{1}<\dots<k_{r}\leq N}{{\operatorname{Sep}_{N}}(k_{1},\dots,k_{r})=r}}\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N}}\right).$
We define
$\tau_{*}^{N}=\sum_{\overset{0<k_{1}<\dots<k_{r}\leq N}{{\operatorname{Sep}_{N}}(k_{1},\dots,k_{r})=r}}\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N}}\right).$
Using $(M1)_{r}$ which has been established, we obtain
$\lim_{N\rightarrow\infty}\frac{\tau_{*}^{N}}{\left(N\sigma_{\rho_{N}}\right)^{r}}=1.$
In addition, since $\mu(A_{*}^{N})\leq\tau_{*}^{N},$ we have
$\limsup_{N\rightarrow\infty}\frac{\mu(A_{*}^{N})}{\left(N\sigma_{\rho_{N}}\right)^{r}}\leq 1.$
(d) We first assume that the point $x$ is not periodic and define
(B.1) $\bar{\tau}_{*}^{N}=\sum_{k_{j}/N\in I_{j}}\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right).$
Recall that
$A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}$ $=$ $\left\{\omega\in\Omega:\,\exists\,k_{j}\,\,\text{such that}\,\,k_{j}/N\in I_{j}\,\,\text{and }\,\,\omega\in\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right\}$
$=$ $\bigcup_{k_{j}/N\in I_{j}}\left(\bigcap_{j=1}^{N}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right).$
Using $(M1)_{r}$ which has been proved, we obtain
$\lim_{N\rightarrow\infty}\frac{\bar{\tau}_{*}^{N}}{N^{r}\Pi_{j=1}^{r}\left(\sigma_{\rho_{N},i_{j}}|I_{j}|\right)}=1.$
Therefore, it suffices to show that
$\lim_{N\rightarrow\infty}\frac{\mu\left(A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}\right)}{\bar{\tau}_{*}^{N}}=1.$
By Boole’s inequality, we have
$\mu\left(A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}\right)\leq\bar{\tau}_{*}^{N}$
and
(B.2) $\mu\left(A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}\right)\geq\bar{\tau}_{*}^{N}-\sum_{t=r+1}^{2r}\bar{\tau}_{t}^{N},$
where
$\bar{\tau}_{t}^{N}=\sum_{\overset{0<k_{1}<\dots<k_{r}\leq N}{{\operatorname{Sep}_{N}}(k_{1},\dots,k_{r})=r}}\mu\left(\Omega^{k_{1}}_{{\rho}_{N}}\cap\dots\cap\Omega^{k_{r}}_{{\rho}_{N}}\cap\Omega^{k_{r+1}}_{{\rho}_{N}}\cap\cdots\cap\Omega^{k_{t}}_{{\rho}_{N}}\right),$
because $\Omega_{\rho_{N},i_{j}}\subset\Omega_{\rho_{N}}=\{y:\phi(y)>C\rho_{N}\}$ for all $i_{j},$ $1\leq j\leq r,$ and some $C>0.$
We prove that
$\frac{\bar{\tau}_{r+1}^{N}}{\bar{\tau}_{*}^{N}}\to 0\,\,\,\text{as}\,\,\,N\to\infty.$
The proof of $\bar{\tau}_{t}^{N}/\bar{\tau}_{*}^{N}\to 0$ for other values of $t$ follows similarly. We consider several cases depending on the size of $p=\min_{j=1,\dots,r}|k_{r+1}-k_{j}|$.
@@
(i) Since $x$ is non-periodic, for each fixed $p_{0},$ we have $\Omega_{{\rho}_{N}}\mathrel{{$\bigcap$}}f^{-p}\Omega_{{\rho}_{N}}=\emptyset$ for all $1\leq p<p_{0},$ provided that $N$ is large enough. Therefore, the contribution from terms with $p<p_{0}$ is zero.
@@
(ii) Next, suppose that $p_{0}\leq p\leq\varepsilon\operatorname{ln}N.$ Specifically, we may assume that $k_{r+1}=k_{\bar{j}}+p$ for some ${\bar{j}}\leq r$ (the case where $k_{r+1}=k_{\bar{j}}-p$ is analogous). Define $\bar{A}_{{\rho}_{N}}^{+}$ as the characteristic function of the set
$B\left(x,\left(\frac{a}{C\rho_{N}-L}\right)^{s}\right).$
Then we have
$\bar{A}^{+}_{\rho_{N}}\geq{\mathbf{1}}_{\Omega_{\rho_{N}}},\,{\mathbb{E}}\bar{A}^{+}_{\rho_{N}}-\mu(\Omega_{\rho_{N}})\leq C\rho_{N}^{-(s+1)}\text{ and }\left\|\bar{A}^{+}_{\rho_{N}}\right\|_{BV}\leq 2\mu(\Omega_{\rho_{N}})+4.$
@@
${\hat{A}}_{j}=\begin{cases}\bar{A}_{{\rho}_{N}}^{+},&\text{if }j\neq{\bar{j}};\\ \bar{A}_{{\rho}_{N}}^{+}\left(\bar{A}_{{\rho}_{N}}^{+}\circ f^{p}\right),&\text{if }j={\bar{j}}.\end{cases}$
Let $\Lambda=\max_{x\in\mathbb{S}}|f^{\prime}(x)|.$ Then, we have
@@
$\leq$ $\left(\prod_{j\neq{\bar{j}}}\left({\mathbb{E}}{\hat{A}}_{j}+C\left\|\bar{A}_{{\rho}_{N}}^{+}\right\|_{BV}\theta^{K\operatorname{ln}N}\right)\right)\left({\mathbb{E}}{\hat{A}}_{\bar{j}}+C\left\|\bar{A}_{{\rho}_{N}}^{+}\right\|_{BV}^{2}\Lambda^{p}\theta^{K\operatorname{ln}N}\right)\leq\prod_{j=1}^{r}\left({\mathbb{E}}{\hat{A}}_{j}+N^{-10}\right)$
@@
if $s(n)=K\operatorname{ln}N$ and $K$ is large enough. Noting that ${\mathbb{E}}{\hat{A}}_{\bar{j}}={\mathbb{E}}\left(\bar{A}_{{\rho}_{N}}^{+}(\bar{A}_{{\rho}_{N}}^{+}\circ f^{p})\right)\leq C\sigma_{{\rho}_{N}}\theta^{p}$ by (2.7), we conclude that each term for $p$ as in the case (ii) is at most $C\sigma_{{\rho}_{N}}^{r}\theta^{p}.$ Summing over all $r$ tuples and over all $p\in[p_{0},\varepsilon\operatorname{ln}N],$ we obtain a total contribution of $CN^{r}\sigma_{\rho_{N}}^{r}\theta^{p_{0}}$ for some constant $C$ independent of $p_{0}.$
@@
(iii) The terms where ${\varepsilon}\operatorname{ln}N<p\leq K\operatorname{ln}N$ contribute $O(\sigma_{\rho_{N}}^{r+\eta})$ for some $\eta=\eta({\varepsilon})>0.$ Therefore, the total contribution of those terms is $O\left(N^{r}(\operatorname{ln}N)\;\sigma_{\rho_{N}}^{r+\eta}\right).$
@@
(iv) The terms where $p>K\operatorname{ln}N$ for sufficiently large $K$ contribute $O(\sigma_{\rho_{N}}^{r+1}),$ similar to the proof of $(M1)_{r}.$ Hence, the total contribution of these terms is $O\left((N\sigma_{\rho_{N}})^{r+1}\right).$
Summing the estimates from cases (i)–(iv), we observe that $\bar{\tau}_{r+1}^{N}$ is negligible in comparison to $\tau_{*}^{N}.$ This completes the proof of $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$ for non-periodic points.
Next, we consider the case where the point $x$ is periodic. Suppose the period of $x$ is $q,$ i.e., $f^{q}(x)=x.$ We use the notation $\tilde{\Omega}_{\rho_{N},j}$ as defined in (2.6) and define
$\tilde{A}_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}=\left\{\omega\in\Omega:\,\exists\,k_{j}\,\,\text{such that}\,\,k_{j}/N\in I_{j}\,\,\text{and }\,\,\omega\in\tilde{\Omega}^{k_{j}}_{{\rho}_{N},i_{j}}\right\}.$
We now proceed to prove $(M4)_{I_{1},\dots I_{r};i_{1},\dots,i_{r}}$ for $\tilde{A}_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}.$
Observe that
$f^{-p_{0}}\Omega_{{\rho}_{N},j}=\left(f^{-(p_{0}-q)}\Omega_{{\rho}_{N},j}\mathrel{{\bigcap}}f^{-p_{0}}\Omega_{{\rho}_{N},j}\right)\mathrel{{\bigcup}}\tilde{\Omega}_{\rho_{N},j}.$
Indeed if $f^{k}y\in\Omega_{{\rho}_{N},j},$ $f^{n}y\in\Omega_{{\rho}_{N},j}$ can only occur if $n=k+iq$ for $i\in\mathbb{Z}.$ Therefore, $(p_{0}-k)$ must be a multiple of $q$. In this case, $f^{p_{0}}y\in\Omega_{{\rho}_{N},j}$ implies $f^{p_{0}-q}y\in\Omega_{{\rho}_{N},j}$, since $f$ is expanding.
In addition, we have
(B.3) $f^{-(p_{0}-q)}\Omega_{{\rho}_{N},j}\mathrel{{$\bigcap$}}f^{-p_{0}}\Omega_{{\rho}_{N},j}=\emptyset$
whenever $0<{\varepsilon}<\left(\gamma^{q/s}-1\right)c_{j}/\left(\gamma^{q/s}+1\right),$ where $\varepsilon$} is from the definition of
$\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\},$
and $\gamma=\min_{x\in\mathbb{S}}|f^{\prime}(x)|>1.$ If $f^{p_{0}-q}(y)\in\Omega_{{\rho}_{N},j},$ we have
$\left|f^{p_{0}}(y)-x\right|\geq\gamma^{q}\left|f^{p_{0}-q}(y)-x\right|\geq\gamma^{q}\left(\frac{a}{\rho_{N}(c_{j}+\varepsilon)+L}\right)^{s}>\left(\frac{a}{\rho_{N}(c_{j}-\varepsilon)-L}\right)^{s}$
where $L=\max_{x\in\mathbb{S}}\left|\psi(x)\right|.$ Therefore, we conclude that $f^{p_{0}}(y)\notin\Omega_{{\rho}_{N},j},$ and (B.3) follow.
Hence $\tilde{\Omega}_{\rho_{N},j}=f^{-p_{0}}\Omega_{{\rho}_{N},j}$ for $0<{\varepsilon}<\frac{(\gamma^{q/s}-1)c_{j}}{\gamma^{q/s}+1}$ and $\mu\left(\tilde{\Omega}_{\rho_{N},j}\right)=\mu\left(f^{-p_{0}}\Omega_{{\rho}_{N},j}\right)=\sigma_{\rho_{N},j}.$
We will show that
$\lim_{N\rightarrow\infty}\frac{\mu\left(\tilde{A}_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}\right)}{N^{r}\Pi_{j=1}^{r}\left(\sigma_{\rho_{N},i_{j}}|I_{j}|\right)}=1.$
Note that
$\sum_{k_{j}/N\in I_{j}}\mu\left(\bigcap_{j=1}^{r}\tilde{\Omega}^{k_{j}}_{{\rho}_{N},i_{j}}\right)=\sum_{k_{j}/N\in I_{j}}\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)=\bar{\tau}_{*}^{N},$
which coincides with the non-periodic case. We prove $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$ using the inequality:
(B.4) $\mu\left(\tilde{A}_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}\right)\geq\bar{\tau}_{*}^{N}-\sum_{t=r+1}^{2r}\tilde{\tau}_{t}^{N},$
where
$\tilde{\tau}_{t}^{N}=\sum_{k_{j}/N\in I_{j}}\mu\left(\tilde{\Omega}^{k_{1}}_{{\rho}_{N},i_{1}}\cap\dots\cap\tilde{\Omega}^{k_{r}}_{{\rho}_{N},i_{r}}\cap\tilde{\Omega}^{k_{r+1}}_{{\rho}_{N},i_{r+1}}\cap\cdots\cap\tilde{\Omega}^{k_{t}}_{{\rho}_{N},i_{t}}\right).$
We now show that $\frac{\tilde{\tau}_{r+1}^{N}}{\bar{\tau}_{*}^{N}}\to 0\,\,\,\text{as}\,\,\,N\to\infty.$ The proof for other values of $t$ follows similarly. We consider several cases depending on the value of $p=\min_{j=1,\dots,r}|k_{r+1}-k_{j}|$.
(1) If $p<p_{0}$ where $p_{0}$ is specified in the definition of $\tilde{\Omega}_{{\rho}_{N},j},$ we take $j_{0}$ such that $k_{r+1}=k_{j_{0}}+p$ (the case where $k_{r+1}=k_{j_{0}}-p$ is similar). Since $d(I_{i},I_{j})>{\varepsilon}$ for $i\neq j,$ both $k_{r+1}/N$ and $k_{j_{0}}/N$ belong to the same component $I_{j},$ thus $i_{r+1}=i_{j_{0}}.$ We then obtain
$\tilde{\Omega}^{k_{j_{0}}}_{{\rho}_{N},i_{j_{0}}}\cap\tilde{\Omega}^{k_{r+1}}_{{\rho}_{N},i_{r+1}}\subset f^{-k_{j_{0}}}\left(f^{-p_{0}}\Omega_{{\rho}_{N},i_{j_{0}}}\mathrel{{\bigcap}}\left(f^{-p}\Omega_{{\rho}_{N},i_{j_{0}}}^{c}\cap\cdots\cap f^{-p-p_{0}+1}\Omega_{{\rho}_{N},i_{j_{0}}}^{c}\right)\right)=\emptyset.$
(2) If $p\geq p_{0},$ we take $C>0$ such that
$\Omega_{\rho_{N},i_{j}}\subset\Omega_{\rho_{N}}=\{y:\phi(y)>C\rho_{N}\}$
for all $i_{j},$ $1\leq j\leq r.$ It follows that
$\tilde{\Omega}_{{\rho}_{N},i_{j}}\subset f^{-p_{0}}\Omega_{{\rho}_{N},i_{j}}\subset f^{-p_{0}}\Omega_{\rho_{N}}$
and hence
$\mu\left(\mathrel{{\bigcap}}_{j=1}^{r+1}\tilde{\Omega}^{k_{j}}_{{\rho}_{N},i_{j}}\right)\leq\mu\left(\mathrel{{\bigcap}}_{j=1}^{r+1}f^{-p_{0}}\Omega^{k_{j}}_{{\rho}_{N}}\right)=\mu\left(\mathrel{{\bigcap}}_{j=1}^{r+1}\Omega^{k_{j}}_{{\rho}_{N}}\right).$
Thus, by the same arguments as (ii)-(iv) in the non-periodic case, we conclude that $\tilde{\tau}_{r+1}^{N}$ is negligible compared to $\bar{\tau}_{*}^{N}.$ This completes the proof of $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$ for periodic points. ∎
@@
Since $H_{N,{\bar{\varepsilon}}}\subset\tilde{H}_{k,{\bar{\varepsilon}}}$ for $2^{k}\leq N\leq 2^{k+1},$ it suffices to show that $\tilde{H}_{k,{\bar{\varepsilon}}}$ occurs only for finitely many $k.$ Using (2.7) and similar method to the proof of Proposition 2.3 we conclude that
@@
Hence $\sum_{k}\mu(\tilde{H}_{k,{\bar{\varepsilon}}})<\infty$ and the result follows from the classical Borel-Cantelli lemma. ∎
@@
(a) Let $\hat{T}_{N,j}=\sum_{n=1}^{N}{\mathbf{1}}_{{\hat{B}}_{N,j}}\circ f^{n},$ where ${\hat{B}}_{N,j}$} in given by (3.8). Since $\frac{{\bar{{\bar{S}}}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\leq\hat{T}_{N,j},$ we have
${\mathbb{E}}\left(\left(\frac{\bar{\bar{S}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\right)^{m}\right)\leq{\mathbb{E}}\hat{T}_{N,j}^{m}.$
Next, we consider
${\mathbb{E}}\left(\binom{\hat{T}_{N,j}}{m}\right)=\sum_{1\leq k_{1}<\cdots<k_{m}\leq N}\mu\left(\bigcap_{j=1}^{m}f^{-k_{j}}{\hat{B}}_{N,j}\right)$
and prove that
(B.5) $\sum_{1\leq k_{1}<\cdots<k_{m}\leq N}\mu\left(\bigcap_{j=1}^{m}f^{-k_{j}}{\hat{B}}_{N,j}\right)\leq C\left(N\mu\left(\hat{B}_{N,j}\right)\right)^{m}$
for non-periodic $x.$ Since $j\leq 0$ and $s<1$, we conclude that (B.5) is less than $\frac{C_{1}}{(\operatorname{ln}N)^{sj\delta m}}$. For $y\in{\hat{B}}_{N,j}$ we have $\phi(y)^{m}\left(N\mu\left(\hat{B}_{N,j}\right)\right)^{m}\leq C_{2}N^{m/s}(\operatorname{ln}N)^{\delta m},$ because on this set
$N^{1/s}(\operatorname{ln}N)^{j\delta}<\phi(y)\leq N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}.$
Thus combining (B.5) with the identity
${\mathbb{E}}\hat{T}_{N,j}^{m}=\sum_{k=0}^{m}\left(\sum_{j=0}^{k}(-1)^{k-j}\binom{k}{j}k^{m}\right){\mathbb{E}}\left(\binom{\hat{T}_{N,j}}{k}\right)$,
we obtain
${\mathbb{E}}{\bar{{\bar{S}}}}_{N,j}^{m}\leq CN^{m/s}(\operatorname{ln}N)^{\delta m}.$
We now prove (B.5). We consider several cases depending on the value of $p=\!\!\!\min_{1\leq j\leq m-1}(k_{j+1}\!-\!k_{j}).$
(i) Since $x$ is non-periodic, for each fixed $p_{0},$ we have
${\hat{B}}_{N,j}\cap f^{-p}{\hat{B}}_{N,j}=\emptyset\,\,\,\text{for all}\,\,\,p<p_{0}$
whenever $N$ is sufficiently large because $x$ is non-periodic. Consequently, the contribution of terms corresponding to tuples with $p\leq p_{0}$ vanishes for large $N.$
@@
(ii) If $p_{0}\leq p\leq\varepsilon\operatorname{ln}N,$ suppose that $\operatorname{Sep}_{N}(k_{1},\dots,k_{m})=l<m.$ Arguing as in case $(M4)_{*}$(ii) of the proof of Proposition 2.3, we conclude that each term is smaller than $C\mu\left(\hat{B}_{N,j}\right)^{l}\theta^{p}$ for fixed $p.$ Since the total number of terms is $O\left(N^{l}p^{m-l}\right),$ it follows that the total contribution of the terms with $p={\hat{p}}$ is bounded above by $CN^{l}{\hat{p}}^{m-l}\mu\left(\hat{B}_{N,j}\right)^{l}\theta^{\hat{p}}$. Summing over tuples with ${\hat{p}}\in[p_{0},{\varepsilon}\operatorname{ln}N],$ we obtain a total contribution of
$CN^{l}{p_{0}}^{m-l}\mu\left(\hat{B}_{N,j}\right)^{l}\theta^{p_{0}}$
for some constant $C$ independent of $N,$ where $p_{0}$ is sufficiently large.
@@
(iii) If $\varepsilon\operatorname{ln}N<p\leq K\operatorname{ln}N$ and $\operatorname{Sep}_{N}(k_{1},\dots,k_{m})=l,$ then arguing similarly to case $(M4)_{*}$(iii) we obtain that
$\mu\left(\bigcap_{j=1}^{m}f^{-k_{j}}{\hat{B}}_{N,j}\right)\leq C\mu\left(\hat{B}_{N,j}\right)^{l+\eta}.$
@@
Summing over all $r$ tuples and over all $p\in(\varepsilon\operatorname{ln}N,K\operatorname{ln}N],$ we conclude that the total contribution is at most $C\left(N\mu\left(\hat{B}_{N,j}\right)\right)^{l}\mu\left(\hat{B}_{N,j}\right)^{\eta}\left(\operatorname{ln}N\right)^{m-l}.$